### v1

# DRS: Deep Question Reformulation With Structured Output

Abstract
Question answering is a fundamental capability of large language models (LLMs). However, when people encounter completely new knowledge texts, they often ask questions that the text cannot answer due to a lack of understanding of the knowledge. Recent research shows that large language models identify the unanswerability of questions, but they lack the ability to help people reformulate their questions. Even powerful models like GPT-3.5 perform poorly in this regard. To enhance the ability of LLMs to assist humans in reformulating questions to extract relevant knowledge from new documents, we propose a zero-shot method called DRS: Deep Question Reformulation With Structured Output. Our proposed method leverages large language models and the DFS-based algorithm to iteratively search for possible entity combinations and constrain the output with certain entities, effectively improving the capabilities of large language models in this area. Extensive experimental results show that our zero-shot DRS method significantly improves the reformulation accuracy of GPT-3.5 from 23.03% to 70.42% and effectively improves the score of open-source large language models, such as Gemma2-9B, from 26.35% to 56.75%.
1 Introduction
Question answering is a fundamental capability of large language models (LLMs). In recent years, as the abilities of these models advance significantly from GPT-3 and InstructGPT to GPT-4 Brown et al. (2020); Ouyang et al. (2022); OpenAI et al. (2024), their performance on various question-answering benchmarks improves markedly. Although large language models demonstrate a strong ability to answer a wide range of questions, human knowledge is inherently limited. When people encounter new, unfamiliar texts in specific domains, they often pose questions for which answers cannot be found in the text itself. These are typically referred to as unanswerable questions. Previous research describes this phenomenon as "presupposition errors" and notes that, in real-world scenarios, over 30% of questions cannot be answered due to this issue Gao et al. (2023a); Yu et al. (2023).
Figure 1: An example of question reformulation using large language models.
Since the advent of BERT and GPT-3 Brown et al. (2020); Devlin et al. (2019), language models have rapidly evolved in question-answering tasks Tan et al. (2023); Wang (2022); Nassiri and Akhloufi (2023); Puri et al. (2020). In recent years, many studies aim to help these models recognize the unanswerability of certain questions, encouraging them to respond appropriately to reduce hallucinations and increase reliability Rajpurkar et al. (2018); Deng et al. (2024); Asai and Choi (2020); Sulem et al. (2021, 2022). However, most of these studies focus solely on making models aware of unanswerable questions, without addressing the modification of unanswerable questions into answerable ones while preserving key entities, which could help users extract valuable information they seek from texts in unfamiliar domains Zhao et al. (2024). More importantly, according to Faustini et al. (2023), in a large-scale industrial experiment, rephrasing unanswerable questions posed to virtual assistants significantly enhances the user experience for millions, which highlights the importance of effectively leveraging LLMs to assist people in question reformulation.
Recent research has begun exploring the ability of large language models to assist in reformulating unanswerable questions. For instance, Zhao et al. (2024) constructed a challenging and high-quality benchmark dataset to evaluate the question-reformulation capabilities of these models. Benchmark tests conducted with well-known models like GPT-3.5 and GPT-4 Ouyang et al. (2022); OpenAI et al. (2024) reveal that, even with the use of Chain of Thought (CoT) reasoning Wei et al. (2023), these models fail to achieve ideal performance in question reformulation. This inspires us to explore effective approaches for leveraging LLMs in question reformulation, enabling people to extract valuable knowledge from unfamiliar texts.
Firstly, to enhance accuracy and efficiency in question reformulation evaluation, we introduce a simple yet effective GPT-4o-mini evaluator as a replacement for the existing Llama2-7B evaluator, which suffers from poor generalization and slow performance. Our GPT-4o-mini evaluator retains the original evaluation criteria while significantly improving accuracy, achieving over 90% compared to the original evaluator’s accuracy of approximately 50%.
Secondly, to enable large language models to effectively assist people in reformulating unanswerable questions from unfamiliar texts, we propose a zero-shot method called DRS: Deep Question Reformulation with Structured Output. This method begins by extracting and filtering entities from the original question, then applies a DFS-based algorithm to identify possible combination of entities, construct text-based statements, and generate corresponding answerable questions - both presented as structured outputs constrained to include selected entities. Finally, the large language model selects the most reliable question from a set of candidates as the final output.
Our method enhances the relevance of reformulated questions to the original by employing a DFS-based algorithm and constraining the model to produce structured outputs, maximizing the likelihood of users obtaining the desired answers. Unlike baseline methods from previous research, our approach takes a more holistic perspective on question generation; the entity constraints ensure greater reliability in the generated questions, resulting in significant accuracy improvements on benchmarks. We conducted extensive evaluations using various LLMs on the dataset proposed by Zhao et al. (2024), which includes texts and questions from diverse sources. The results show that our method significantly boosts various large language models’ reformulation accuracy across all datasets, clearly demonstrating its effectiveness.
In summary, our contributions in this paper are as follows:
(i) We propose a zero-shot method called DRS, which enables LLMs to effectively reformulate unanswerable questions by using a DFS-based combination search to generate structured and reliable outputs.
(ii) Extensive experiments show that zero-shot DRS significantly improves LLMs’ performance, surpassing few-shot CoT by over 100%.
(iii) We introduce the high-speed, high-accuracy GPT-4o-mini evaluator to replace the slow, low-performing Llama2-7B evaluator, enhancing fairness and effectiveness.
Figure 2: The process of our zero-shot DRS method, which mainly contains three different part - entity extraction, dfs combination search with question generation, final candidate selection.
2 Related Work
Question answering has long been a core focus in natural language processing (NLP) Wang (2022). For many years, researchers have developed new datasets to explore the capabilities of language models in this area, with well-known examples such as CBT, SearchQA, and others Hill et al. (2016); Dunn et al. (2017). These datasets primarily assess a model’s ability to answer questions, without considering the possibility that some questions may be unanswerable.
Since the advent of GPT-3 Brown et al. (2020), the rapid development of large language models has increasingly focused attention on unanswerable questions, whether due to ambiguity or limitations in knowledge Deng et al. (2024); Gao et al. (2023b); Yin et al. (2023). Many studies have proposed methods to address question ambiguities for LLMs, often targeting issues like missing qualifiers that can lead to varied answers Kim et al. (2024); Zhang and Choi (2023).
Extensive research on questions outside the knowledge scope of LLMs has focused on guiding models toward self-reflection, helping them recognize when they lack sufficient knowledge to answer a question and thus reducing hallucinations Yin et al. (2023); Ji et al. (2023); Tonmoy et al. (2024). However, the research often overlooks a common real-life scenario: when people encounter unfamiliar knowledge domains and texts, they may pose questions that appear relevant but are ultimately unanswerable Zhao et al. (2024).
The issue of unanswerable questions in QA systems was first formally addressed in a dataset developed by Rajpurkar et al. (2018). This pioneering work has since been expanded by various studies, such as Yu et al. (2023) analyzing Google search queries and Kim et al. (2023) examining Reddit discussions. Both teams focused on identifying questions with incorrect assumptions or presuppositions that made them impossible to answer accurately based solely on the given document.
Recently, Zhao et al. (2024) advanced research in this field by creating a high-quality dataset of document-related unanswerable questions to explore how LLMs can help people reformulate such questions based on the document. They applied basic approaches, including zero-shot, few-shot, and CoT prompting, to several prominent large language models to evaluate their effectiveness in reformulating unanswerable questions. However, the results were suboptimal. Therefore, in this paper, we propose the zero-shot DRS method to significantly improve large language models’ ability to assist users in question reformulation, further unlocking their potential in this area.
3 Methodology
In this paper, we propose a new zero-shot method called DRS: Deep Question Reformulation with Structured Output, which significantly improves LLMs’ ability to reformulate unanswerable questions based on a given document, helping users obtain desired answers in unfamiliar knowledge domains.
Our approach consists of three main steps: (i) Entity extraction and filtering, where large language models are used to extract core entities. (ii) Entity combination search and structured output generation, which employs a DFS-based algorithm to identify entity combinations to form answerable questions. (iii) DFS pruning and candidate selection, where we limit the number of candidate questions and search depth through DFS pruning, then select the most reliable question from the generated candidates for final output. The detailed process of our zero-shot DRS method is illustrated in Figure 2.
3.1 Entity Extraction and Filtering
When people encounter a text with unfamiliar knowledge, they usually raise questions about the key entities in the text, indicating their desire to understand more about these entities. Therefore, when reconstructing questions, we focus on the key entities from the original question. This ensures that the reconstructed question addresses the content people genuinely care about, rather than an arbitrary, overly simple new question.
For example, when people ask a question like, "When does Rainer Hertrich, the German co-head of EADS, step down?", the important entities are "Rainer Hertrich", "German", and "EADS". These terms represent the key subject of the sentence and the relevant modifiers. Modern large language models can identify important entities from a question. However, these models may also include phrases like "step down", which is a verb phrase, not an important entity, as part of the important entities. This results in the model outputting less precise entities, leading to greater errors in subsequent steps.
Therefore, this process involves two sub-steps:
(i) We use a simple and direct zero-shot prompt to have the large language model extract all entities it considers important from the original question. The goal is to minimize the omission of any potentially important entities.
(ii) We then apply the large language model again with a zero-shot prompt as an effective classifier to classify the previously extracted entities into five categories: subject, object, predicate, attribute, and others. We retain all entities classified as subject, object, and attribute, as these are the most important components of the question, and discard the others.
[…]
Table 1: The detail information of all experimental data for evaluating the performance on the question reformulation task.
Evaluator QA2 BanditQA BBC Reddit Yelp SQuADv2 Average 200 / 247 300 / 507 50 / 59 100 / 113 50 / 51 300 / 736 1000 / 1713 Llama2-7B Zhao et al. (2024) $39.0039.00$ $68.6768.67$ $14.0014.00$ $79.0079.00$ $44.0044.00$ $72.0072.00$ $52.7852.78$ GPT-4o-mini $90.5090.50$ $92.3392.33$ $88.0088.00$ $92.0092.00$ $90.0090.00$ $91.3391.33$ $90.7090.70$
Table 2: The comparison of prediction accuracy between previous Llama2-7B evaluator proposed by Zhao et al. (2024) and our proposed GPT-4o-mini evaluator.
3.2 DFS Combination Search and Structured Question Generation
After obtaining reliable entities, we proceed with question reformulation. We use the DFS algorithm to explore combinations of different entities. For any combination that could reformulate a new question, we allow it to generate a completely new question and store it for later selection. Specifically, in this step, we focus on the following processes:
(1) We use the DFS algorithm to select possible entity combinations. When the number of entities in a combination exceeds half of the filtered entities, we move on to step (2).
(2) We prompt the large language model with a concise zero-shot instruction to generate a structured statement containing all selected entities, based on the chosen entities and corresponding text. This statement serves as a conclusion that can be drawn from the document.
(3) After generating the statement, we ask the large language model to create a structured question that includes all the selected entities and can be answered, based on the generated statement.
(4) We then return the generated question to the large language model and ask it to verify whether the question includes all the necessary entities. If it does, we proceed to the next step (5); if not, we return to step (1) and select a new combination.
(5) If the question contains all the entities, we input it again into the large language model to check whether it can be answered. If it can, we store both the question and the count of its entities for final selection. If not, we discard it and return to step (1) to search for another combination. (Note: When the number of valid questions reaches the set threshold, no further combinations will be attempted.)
In this step, our algorithm uses a DFS search approach to identify possible entity combinations that can be reformulated into new questions. During the generation process, we find that directly prompting the large language model to reshape the question often results in a low success rate due to the model’s potentially unlimited output content. To address this, we guide the large language model to generate structured outputs by restricting the content required in the generated text. We ensure the reliability of the reformulated question by first generating a statement, then using that statement to create the question. Additionally, to enhance the algorithm’s efficiency, we verify the entities immediately after generating the new question, preventing unnecessary subsequent steps. We also apply pruning to the DFS algorithm by limiting the total number of preselected questions and controlling the search depth.
3.3 Candidate Question Evaluation
In the previous step, we generate a set of new questions through search and recorded the number of entities contained in each. Now, we input these questions along with their respective entity counts into the large language model for reassessment of their answerability. Using a straightforward prompt, we guide the model to evaluate and select the optimal question. When the model determines that multiple questions are answerable based on the document, it selects the question with the highest number of entities and outputs its corresponding label, yielding the final reformulated question.
4 Experiments
4.1 Datasets
In this paper, we utilize the newly constructed high-quality dataset CouldAsk [^1] https://huggingface.co/datasets/wentingzhao/couldask, introduced by Zhao et al. (2024). This dataset comprises subsets from six diverse sources, including Yelp, BBC, SQuAD, Reddit, and others, featuring over 4,000 high-difficulty, high-quality data points filtered by both human annotators and GPT-4. For our study, we focus on over 1,700 unanswerable questions selected from all subsets as test data for model reformulation. The detailed information about our experimental data is provided in Table 1.
4.2 Large Language Models
To effectively evaluate our methodology and compare it with baseline approaches from previous studies, we conduct all experiments using four different LLMs.
Closed-Source Models. To balance experimental cost and effectiveness, we primarily use two widely adopted models from OpenAI: GPT-3.5 and GPT-4o-mini.
Open-Source Models. We also perform experiments using two well-known models developed by Google and Alibaba: Gemma2-9B  [^2] https://huggingface.co/google/gemma-2-9b-it Team et al. (2024) and Qwen2.5-7B  [^3] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct Yang et al. (2024); Team (2024).
4.3 Metric
We employ the metric Accuracy, in line with previous research, to measure the proportion of unanswerable questions successfully reformulated. This evaluation involves two criteria:
(i) We use GPT-4o-mini as an evaluator to generate reasoning steps and assess whether the reformulated question can be answered using the relevant text. (ii) We also use GPT-4o-mini as an entity detector to measure the overlap of entities between the reformulated question and the original question. If the number of overlapping entities exceeds half of the labeled entities in the dataset, we consider the reformulated question likely to help the inquirer obtain the desired information.
A reformulated question is deemed successful and counted towards accuracy only if it satisfies both criteria.
4.4 Implementation Details
In our main experiments using the proposed DRS method, we set the temperature to 0.0 for all large language models to ensure reproducibility. Similarly, for the GPT-4o-mini evaluator, we set the temperature to 0.0 to maintain reliability and consistency in the model’s judgments.
For the two GPT series models, we use the official OpenAI API [^4] https://openai.com/ for inference. For the two open-source models, we deploy them on a single A6000 GPU using bfloat16 precision and use the generate function for text generation. (Additional details are provided in Appendix A).
Model Method    QA2 BanditQA     BBC Reddit     Yelp SQuADv2 Average GPT-3.5 Zero-Shot $28.7428.74$ $11.4111.41$ $13.5613.56$ $18.5818.58$ $19.6119.61$ $36.6936.69$ $21.4321.43$ Zero-Shot w/ CoT $29.1529.15$ $15.6315.63$ $13.5613.56$ $20.3520.35$ $15.6915.69$ $43.7943.79$ $23.0323.03$ Few-Shot $32.7932.79$ $22.6922.69$ $15.2515.25$ $23.0123.01$ $29.4129.41$ $38.4638.46$ $26.9426.94$ Few-Shot w/ CoT $44.9444.94$ $48.7848.78$ $16.9516.95$ $18.5818.58$ $25.4925.49$ $41.2241.22$ $32.6632.66$ Zero-Shot DRS (ours) 81.80 73.20 62.71 75.22 66.67 62.90 70.42 GPT-4o-mini Zero-Shot $36.4436.44$ $15.0815.08$ $15.2515.25$ $23.0123.01$ $15.6915.69$ $47.9347.93$ $25.5725.57$ Zero-Shot w/ CoT $49.3949.39$ $37.5037.50$ $42.3742.37$ $17.7017.70$ $35.2935.29$ $50.1050.10$ $38.7338.73$ Few-Shot $36.4436.44$ $17.8017.80$ $16.9516.95$ $21.2421.24$ $21.5721.57$ $46.5546.55$ $26.7626.76$ Few-Shot w/ CoT $61.1361.13$ $52.8552.85$ $47.4647.46$ $43.3643.36$ $56.8656.86$ $59.5759.57$ $53.5453.54$ Zero-Shot DRS (ours) 88.26 80.16 79.66 83.19 78.43 78.30 81.33 Qwen2.5-7B Zero-Shot $46.1546.15$ $24.0524.05$ $25.4225.42$ $23.8923.89$ $25.4925.49$ $46.7046.70$ $31.9531.95$ Zero-Shot w/ CoT $47.1847.18$ $44.2144.21$ $27.1227.12$ $35.4035.40$ $31.3731.37$ $53.5553.55$ $39.8139.81$ Few-Shot $39.6839.68$ $25.1425.14$ $15.2515.25$ $23.8923.89$ $19.6119.61$ $39.6439.64$ $27.2027.20$ Few-Shot w/ CoT $54.2854.28$ $45.5945.59$ $25.4225.42$ $25.6625.66$ $33.3333.33$ $50.6450.64$ $39.1539.15$ Zero-Shot DRS (ours) 77.73 69.84 55.93 53.10 58.82 60.55 62.66 Gemma2-9B Zero-Shot $36.0336.03$ $21.3321.33$ $20.3420.34$ $18.5818.58$ $19.6119.61$ $42.2142.21$ $26.3526.35$ Zero-Shot w/ CoT $29.1529.15$ $17.1217.12$ $22.0322.03$ $23.0123.01$ $19.6119.61$ $32.9432.94$ $23.9623.96$ Few-Shot $26.3226.32$ $16.1716.17$ $25.4225.42$ $15.0415.04$ $17.6517.65$ $29.3929.39$ $21.6721.67$ Few-Shot w/ CoT $22.6722.67$ $18.0718.07$ $10.1710.17$ $24.7824.78$ $25.4925.49$ $27.4227.42$ $21.4321.43$ Zero-Shot DRS (ours) 59.92 60.73 55.93 59.29 49.02 55.62 56.75
Table 3: The main experimental results of four different large language models on full data across six different datasets, where best results are highlighted in bold font.
5 GPT Evaluator
In the previous paper, Zhao et al. (2024) fine-tuned the Llama2-7B model Touvron et al. (2023) on the training dataset for the sequence classification task to determine whether a question could be answered from the corresponding text. This approach achieved high scores on their validation dataset. However, in our experiments, we find that its generalization performance is suboptimal; even on the CouldAsk dataset, which includes training data, it still yields extremely poor classification results.
Given the situation described above, we find it necessary to propose a reliable and efficient evaluation method. Therefore, in this paper, we choose the GPT-4o-mini model, which demonstrates excellent language understanding capabilities, as the evaluation model, ensuring both strong evaluation performance and cost-effectiveness.
To demonstrate the strong capabilities of the GPT-4o-mini model, we select a subset of labeled questions from each portion of the CouldAsk dataset. We then have both evaluation models perform classification predictions and record their classification accuracy on this data, with the results shown in Table 2.
We observe that GPT-4o-mini consistently achieves high scores across all tests, whereas the fine-tuned Llama2-7B model shows significantly lower accuracy than GPT-4o-mini and exhibits substantial score variation across different datasets (ranging from a high of 79% to a low of 14%). For the average accuracy across six different datasets, our GPT-4o-mini evaluator achieves over 90%, while the Llama2-7B evaluator only reaches 52%, performing slightly better than random predictions. This clearly indicates that the fine-tuned model lacks generalizability across diverse data. Therefore, using GPT-4o-mini as the evaluator is not only more reliable but also provides faster evaluation and lower operational costs.
Figure 3: Model accuracy across four large language models with varying numbers of candidate questions, evaluated on six datasets and their average.
6 Experimental Results
To demonstrate the effectiveness of our proposed zero-shot DRS method, we test it on six datasets, totaling over 1,700 data points, using four different large language models. We conduct a fair and thorough comparison between our method and four baseline methods from previous research, with all accuracy scores presented in Table 3. Additionally, we adjust relevant parameters, such as temperature and the number of candidate questions, in the zero-shot DRS method and conduct multiple experiments. The results are shown in Figure 4 and Figure 3. Finally, considering GPU resources and API costs, we select a subset of test data for additional experiments with more models, as listed in Appendix B.
Effective DRS. Table 3 clearly demonstrates the strong capabilities of our proposed zero-shot DRS method. Our method significantly outperforms all baselines across all datasets and models. Using the GPT-3.5 model, the average accuracy in the zero-shot setting increases from 23% to 70%, nearly tripling the original performance — a remarkable improvement that underscores the effectiveness of our approach. With the Gemma2-9B model, while the results are lower than those of GPT-3.5, our method still boosts the zero-shot accuracy from 26% to 57%, achieving an improvement of over 100%. Moreover, even when compared to few-shot baselines, our zero-shot method consistently outperforms them. On GPT-3.5, the 70% average accuracy is more than double the 33% accuracy of the few-shot CoT method. For Gemma2-9B, the few-shot accuracy is even lower than the zero-shot baseline.
In addition to the significant improvements observed with Gemma2-9B and GPT-3.5, both GPT-4o-mini and Qwen2.5-7B also show remarkable enhancements. For GPT-4o-mini, the accuracy achieved by zero-shot DRS is more than double that of zero-shot CoT, and still shows a 52% improvement over the few-shot CoT results. While Qwen2.5-7B exhibits the smallest improvement ratio among the four models, it still achieves over a 57% enhancement compared to both zero-shot and few-shot CoT. These comparisons collectively demonstrate the effectiveness of the zero-shot DRS method in question reformulation.
Robust DRS. In the process of our DRS method, several steps involve using large language models to generate corresponding answers. The model’s temperature influences the output, affecting the diversity and reliability of the responses. Due to the continuity and interrelation of these steps, errors can propagate and accumulate. Therefore, we apply different temperatures to various large language models in our method, and the experimental results are presented in Figure 4.
We observe that the average scores remain remarkably stable as the temperature increases. The score differences between different temperature settings are at most within 3 percentage points, which is negligible compared to the performance improvement brought by our method. Additionally, we notice that GPT-3.5 and GPT-4o-mini achieve better scores when the temperature is set to 0.0 or 0.7. This is likely because, at a temperature of 0.0, the models can more strictly follow the instructions containing entities, while at 0.7, the diversity of the outputs is effectively utilized, providing more varied options for candidate questions and increasing the likelihood of selecting a better final question. In contrast, the accuracy of Qwen2.5-7B and Gemma2-9B slightly decreases with higher temperatures, as smaller models struggle to follow instructions effectively as the temperature increases.
Figure 4: Model accuracy across four large language models with five different temperature settings, evaluated on six datasets and shows their average.
Impact of the Number of Candidate Questions. In our DRS method, before returning the final output question, we guide the model to generate a certain number of candidate questions. To explore the impact of this parameter on the performance of our method, we conduct experiments with varying numbers of candidate questions and observe how accuracy changes. We set the number of candidate questions from 1 to 5 for each large language model and test on all datasets, with the results shown in Figure 3.
We observe that the number of candidate questions does affect the final accuracy. Notably, when only one candidate question is generated, the accuracy of question reformulation across various models drops more significantly compared to other settings. For instance, with GPT-3.5, setting the number of candidate questions to 1 results in a decrease in accuracy from 70.42% to 63.57%, a drop of nearly 10%. Similarly, GPT-4o-mini ’s accuracy decreases from 81.33% to 78.56%. This is because, when only one question is generated, inaccuracies in entity extraction, coupled with a shallow search depth in our DRS algorithm, make it challenging to ensure significant overlap between the entities in the reformulated question and the true label entities. Consequently, the reformulated question often fails to meet the entity overlap requirement.
However, when the number of candidate questions is set from 2 to 5, the accuracy differences are relatively minor, typically within 3 percentage points (only Qwen2.5-7B shows a large decrease when candidate number is 4), following a general trend of initially increasing and then decreasing. Besides, the curve shows that setting the number of candidate questions to 2 or 3 yields the most ideal results, achieving high accuracy while significantly saving search time. Even more convincing is the fact that, even when the number of candidate questions is set to the least ideal value of 1, the scores obtained by our DRS method still far outperform all other baselines. This strongly demonstrates the power and remarkable robustness of our method.
Model Method Inference Time GPT-3.5 Zero Shot w/ CoT 6.80s Few Shot w/ CoT 7.32s DRS (candidate num = 2) 10.07s DRS (candidate num = 3) 11.44s Gemma2-9B Zero Shot w/ CoT 12.51s Few Shot w/ CoT 13.82s DRS (candidate num = 2) 17.36s DRS (candidate num = 3) 19.65s
Table 4: Average inference time per sample across six datasets for three types of methods.
Inference Time of DRS. In Table 4, we present the average runtime of GPT-3.5 and Gemma2-9B under different methods. For our DRS method, we include the runtime when the number of candidate questions is set to 2 or 3, as Figure 3 shows that optimal performance can already be achieved with 2 or 3 candidates. From Table 4, we observe that the DRS method, implemented based on a DFS search approach, results in longer runtimes compared to the few-shot CoT method, despite the use of effective pruning techniques. Specifically, when the number of candidate questions is 2, the runtime increases by 37% on GPT-3.5 and 25% on Gemma2-9B.
However, despite the slight increase in inference time, the accuracy improvements of the DRS method over the few-shot CoT method are remarkable, with increases of 111% and 160% on the two large language models, respectively. This clearly shows a trade-off between runtime and accuracy, but the accuracy gains of the DRS method far outweigh the impact of the longer runtime. Therefore, we have strong evidence to support that the DRS method is a reasonable and efficient approach.
Figure 5: Comparison of our zero-shot DRS method with four baseline approaches, where our method successfully reformulates the question while all other baselines do not.
7 Case Study
To highlight the advantages of our proposed zero-shot DRS method, we select some examples to demonstrate its effectiveness compared to other baselines, as shown in Figure 5. We observe that the relevant documents primarily discuss mathematical theorems, and the user’s question involves five entities but cannot be answered using the corresponding documents. In these examples, the reformulated questions generated using zero-shot and zero-shot CoT are identical. These questions overlap with the original question on three entities: "conjecture", "integer n", and "primes". However, they still cannot be answered based on the provided documents, making such reformulations unsuccessful.
In contrast, questions generated using the few-shot or few-shot CoT methods are more diverse. However, despite being answerable based on the document, they fail to pass the entity overlap check. Specifically, the questions generated by the few-shot method contain no entities from the original question, while those generated by the few-shot CoT method include only two entities. Both fail to meet the required entity overlap ratio of 0.5 or higher, making these reformulations unsatisfactory.
In comparison, the reformulated question generated by our proposed DRS method satisfies both strict conditions: answerability and entity overlap ratio. The DRS method generates questions through a DFS search process, where the model is guided by the constraints of three entities: "conjecture", "integer n", and "primes". In this process, we first prompt the model to construct a statement based on the document, using these entities. We then use this statement to generate a reformulated question containing the same entities. This approach ensures both answerability and entity overlap, optimizing the model’s ability to help users obtain relevant answers.
From these observations, we conclude that even advanced methods like few-shot or few-shot CoT struggle to produce satisfactory reformulations due to the lack of constraints on generated content. This lack of control often results in outputs that either overly focus on preserving the original question’s entities, leading to unanswerable questions, or prioritize answerability at the expense of entity overlap. In contrast, the DRS method effectively balances these two requirements by systematically searching for entity combinations and generating structured statements and reformulated questions containing the specified entities. This approach unlocks the full potential of large language models, enabling more effective question rephrasing for users.
8 Conclusion
In our paper, we propose a zero-shot DRS method that significantly enhances the performance of large language models on the unanswerable question reformulation task. Our method outperforms all other baselines, including the few-shot CoT approach, across six different datasets and various large language models. We also conduct extensive experiments on key parameters, such as temperature and the number of candidate questions, with results demonstrating the strong capabilities and robustness of our approach. Furthermore, our findings reveal that the previously proposed Llama2-7B evaluator lacks the capability to fairly assess the reformulated questions. We propose using GPT-4o-mini as a more reliable, accurate, and low-latency evaluator for future research. Looking ahead, we aim to develop even more effective methods to harness the full potential of large language models, helping them assist people in understanding unfamiliar documents more effectively.
Limitation
In this paper, our proposed zero-shot DRS method significantly enhances the ability of large language models to reformulate unanswerable questions. However, there are still the following limitations: (i) Despite the use of efficient pruning with the DFS algorithm, the process still requires multiple passes through the document, which increases computational costs. Future research could explore methods to enable large language models to complete the reformulation in a single attempt. (ii) While we conduct tests on six datasets, they do not include documents from specific disciplines, which are common in real-world applications. Future work could focus on developing datasets that cover a broader range of domains, allowing for a more comprehensive evaluation of algorithmic and large language model capabilities.
Ethics Statement
Ethical considerations are of utmost importance in our research endeavors. In this paper, we strictly adhere to ethical principles by exclusively utilizing open-source datasets and employing various models that are either open-source or widely recognized in the scientific community. Our proposed methodology aims to enhance the model’s ability to reformulate unanswerable questions when encountering documents from new knowledge domains in real-world scenarios. We are committed to upholding ethical standards throughout the research process, prioritizing transparency, and promoting the responsible use of technology for the betterment of society.
Acknowledgements
The authors would like to thank the anonymous reviewers for their discussion and feedback.
Appendix A Implementation Details
We use the gpt-3.5-turbo-0125 version for GPT-3.5, gpt-4-0125-preview for GPT-4, and gpt-4o-mini for GPT-4o-mini. All experiments were conducted between November 9th and 25th, 2024.
To ensure fair experimental comparisons and re-evaluate the baseline scores reported in the previous paper, we use the open-source GitHub code provided by Zhao et al. (2024), maintaining consistency in the prompts and LLM parameters. The versions of all OpenAI models also remain consistent with the ones mentioned above.
Appendix B Additional Experiments
To address the constraints of GPU resources and API costs, we randomly select appropriately sized subsets from each test dataset for experiments with three well-known models: GPT-4, Llama3.1-70B, and Gemma2-27B, which demonstrates the generalizability of our approach. We use the complete datasets for Yelp, BBC, and Reddit, while for the three larger datasets, we randomly select 150 samples for the experiments. For comparison, we focus on the results of our DRS method against two representative baselines, zero-shot and few-shot, due to the high computational cost of the models. The experimental results appear in Table 5.
Based on the experimental results, our proposed DRS method demonstrates significant improvements over both zero-shot and few-shot baselines across three newly tested large language models. For instance, on Llama3.1-70B, zero-shot DRS achieves over a 100% improvement in accuracy compared to the baselines. Similarly, on GPT-4, the improvement approaches 100%. Furthermore, with the DRS method, the average reformulation success rate for these two models reaches nearly 70%, which is remarkable. In contrast, Gemma2-27B achieves an average accuracy of only about 50% with DRS, lagging behind the other two models. This discrepancy may be attributed to differences in model size and instruction adherence. Nevertheless, on Gemma2-27B, DRS still boosts accuracy by approximately 125% over the zero-shot baseline and by around 275% over the few-shot baseline. Overall, these results further validate the effectiveness of our DRS method, highlighting its strong generalization and robustness across diverse large language models.
Model Method    QA2 BanditQA     BBC Reddit     Yelp SQuADv2 Average GPT-4 Zero-Shot $44.0044.00$ $30.0030.00$ $28.8128.81$ $25.6625.66$ $21.5721.57$ $64.0064.00$ $35.6735.67$ Few-Shot $48.0048.00$ $29.3329.33$ $37.2937.29$ $23.0123.01$ $35.2935.29$ $60.6760.67$ $38.9338.93$ Zero-Shot DRS (ours) 78.67 72.67 71.19 66.37 62.75 70.00 70.28 Llama3.1-70B Zero-Shot $36.0036.00$ $31.3331.33$ $33.9033.90$ $19.4719.47$ $21.5821.58$ $58.0058.00$ $33.3833.38$ Few-Shot $29.3329.33$ $29.3329.33$ $27.1227.12$ $18.5818.58$ $25.4925.49$ $36.6736.67$ $27.7527.75$ Zero-Shot DRS (ours) 65.33 74.67 67.80 61.95 74.51 66.67 68.48 Gemma2-27B Zero-Shot $10.0010.00$ $12.0012.00$ $23.7323.73$ $18.5818.58$ $17.6517.65$ $42.6742.67$ $20.7720.77$ Few-Shot $28.6728.67$ $8.008.00$ $6.786.78$ $4.424.42$ $13.7313.73$ $20.6720.67$ $13.7113.71$ Zero-Shot DRS (ours) 42.00 64.00 52.54 46.02 37.25 55.33 49.52
Table 5: The additional experimental results of three different large language models on full data across six different datasets, where best results are highlighted in bold font.
Appendix C DRS Prompts
Entity Extraction Prompt Find out all entities in the following question: {question}. The entities you find must be exactly exist in the given question. You should only return the entities, separated by comma and space.
Figure 6: Entity Extraction Prompt
Entity Categorize Prompt Here is a question: {question}. Here is an entity in this question: {entity}. Tell me which category does this entity belong to in the given question - subject, object, predicate, attribute or others. You should only consider the situation of given entity in this specific question. Give your analysis within <analysis> tags, and only return its category name within <answer> tags.
Figure 7: Entity Classification Prompt
Statement and Question Generation Prompt According to the following text: {context}. Generate a statement you could get from the given text that contains all following entities: {entities}. Then you are required to generate a question contains all given entities, which could be answered from your statement. Return the statement within <statement> tags and the question within <question> tags.
Figure 8: Statement and Question Generation Prompt
Entity Overlap Evaluation Prompt Here is an original question: {question}, it contains the following entities: {entities}. Here is a new question: {new_question}. Tell me the number of overlapping entities between the new question and the original question, they do not need to be strictly the same, as long as mentioned, uppercase or lowercase doesn’t matter. Give your analysis within <analysis> tag, and only return the math number of overlap entities within <answer> tags.
Figure 9: Entity Overlap Evaluation Prompt
Question Answerability Evaluation Prompt Here is a long context: {context}. Here is a question: {question}. Tell me whether this question is answerable only according to the information in the provided context. Think carefully and give you analysis within <analysis> tags, then return only the final answer ’yes’ or ’no’ within <answer> tags.
Figure 10: Question Answerability Evaluation Prompt